OpenStudy (anonymous):

If g(x) = x^2 + 3, find g(4). 11 19 16 8

OpenStudy (sumeer):

4^2 + 3 = 16+3= 19

OpenStudy (anonymous):

Was going to ask you to explain but...guess you left :/ thx anyway

OpenStudy (sumeer):

here x is taken as 4 in the function of g(x) . so that you just need to plug 4 to wherever you see x in the function.